Midyear on July 31st, the Chester Corporation's balance sheet reported: Total Assets of $82.572 million Total Common Stock of $2.540 million Cash of $4.020 million Retained Earnings of $17.730 million. What were the Chester Corporation's total liabilities?

Select: 1

$58.282 million.

$64.842 million.

$62.302 million.

$66.322 million.
Answer #1

Total equity=Common stock+Retained earnings

=(2.54+17.73)=$20.27 million

Total assets=Total liabilities+Total equity

Hence Total liabilities=82.572-20.27 million

=$62.302 million